I need help solving this equation.
The attendance at a baseball game was 400 people. Student tickets cost $2 and adult tickets cost $3. total ticket sales were $1050. How many tickets of each type were sold? I need to write two equations to solve this problem.
I got this far:
s= students = $2
a= adults = $3
attendance = 400
tickets sales = $1050

Answer by scott8148(6628)

"The attendance at a baseball game was 400 people" __ s+a=400

" Student tickets cost $2 and adult tickets cost $3. total ticket sales were $1050"
__ s(2)+a(3)=1050 __ 2s+3a=1050